# Estimating the Reynolds number dependence of the chaotic attractor dimension in two-dimensional Kolmogorov flow

### v1

@@ IV.2 Attractor dimension
Here we present the main result of this study – the scaling with $Re$ of the estimated dimension of the chaotic attractor. Implicit rank minimization drives the rank of the embedding space to an approximately minimal rank solution , which we use as (an upper bound for) our numerically estimated dimension. To estimate this minimal rank, we perform a singular value decomposition (SVD) on the covariance of the latent data matrix $\mathbf{Z}$ [26]. The SVD of $\mathbf{Z}\mathbf{Z}^{T}=\mathbf{U}\bm{\Sigma}\mathbf{U}^{T}$ computes the singular vectors $\mathbf{u}_{j}\in\mathbb{R}^{d_{z}}$ where $\mathbf{U}=[\mathbf{u}_{0}\dots\mathbf{u}_{d_{z-1}}]$ (the left singular vectors of $\mathbf{Z}$) and singular values $\sigma_{j}\in\mathbb{R}$ where $\bm{\Sigma}=\text{diag}(\sigma_{0},\dots,\sigma_{d_{z-1}})$ (the square of the singular values of $\mathbf{Z}$). The number of non-negligible singular values in this spectrum is the estimate for $d_{\mathcal{A}}$ in the fully symmetry reduced subspace. Each of these singular vectors represents some important structure of the original data in the latent space, which will be investigated in more detail in §V.
Figure 7: Singular values $\sigma_{i}$ in the singular value decompositions of the covariance of the latent data matrix at $Re=$ 25 (blue), 30 (orange), 35 (green), 40 (red), 60 (purple), 80 (pink), 100 (gray), 200 (brown) and 400 (yellow).
The singular values for each $Re$ considered are reported in figure 7. In all cases there is a sharp drop before $\sigma_{j}$ becomes negligibly small at a critical index $j=c$, corresponding to the rank of the learned, fully symmetry-reduced, latent space. To compute the estimate for $d_{\mathcal{A}}$ in full state space, which we denote by $\tilde{d}_{\mathcal{A}}$, we must add the continuous and discrete symmetries back into the estimate, such that $\tilde{d}_{\mathcal{A}}=c+1+16$. We computed two estimates for $\tilde{d}_{\mathcal{A}}$ at each $Re$ by training two randomly initialized autoencoders from scratch. These estimates were found to be robust to the training – they differed only slightly between models at all $Re$. The smallest estimate at each $Re$ is reported in figure 8.
The non-dimensional Grashof number $G$ in the rigorous upper bound [7, 8] of the global attractor dimension is related to the present definition of $Re$, with monochromatic $n=4$ Kolmogorov forcing, by $G=2^{5/2}\pi^{3}Re^{2}$. This upper bound then scales with $Re$ like $bRe^{4/3}(1+\log{(2^{5/2}\pi^{3}Re^{2})})^{1/3}$, and is indicated by the bold line in figure 8 ($b$ is an unspecified constant). This bound and the lower bound in [9] both scale with the number of degrees of freedom in a two-dimensional domain ($Re^{4/3}$ is proportional to the square of the domain size $L$ over the Kraichnan scale) with a logarithmic correction.
In contrast, our data-driven estimate for the dimensionality of the chaotic attractor explored by long turbulent trajectories exhibits a much weaker scaling at high $Re$. While we observe some non-monotonic behavior in figure 8 at lower $Re$, there is a sharp transition beyond $Re=40$ beyond which an approximate scaling $d_{\mathcal{A}}\sim Re^{1/3}$ is observed – identified with the dotted line in the figure.
@@
Figure 8: The estimated upper bound for the dimension of the turbulent attractor $\tilde{d}_{\mathcal{A}}$ against $Re$. The bold line shows the rigorous upper bound scaling $\sim Re^{4/3}(1+\log{(2^{5/2}\pi^{3}Re^{2})})^{1/3}$ and the dotted line shows the scaling $\sim Re^{1/3}$.
@@
One caveat with our estimates is that our models do not perfectly reconstruct their inputs, and this error does increase with $Re$. However, the markedly weaker scaling with $Re$ is consistent with the formation of coherent structures which are associated with a subset of wavenumber pathways in a Fourier representation. We now explore the representation of the turbulence learned in these models by decoding individual components from the SVD.
@@ V.1 Modal structures
We first examine the contribution to the vorticity field reconstruction from each of the singular vectors (modes) in the SVD of $\mathbf{Z}\mathbf{Z}^{T}$. To do this, we project low and high dissipation snapshots onto the $\ell=0$ mode and combinations of this mode with higher modes. The projection operator onto mode $\ell$ in the embedding space for some vorticity field $\omega$ is defined by
@@
Figure 9: The contributions $\mathcal{D}(\mathcal{P}_{0}(\omega)\mathbf{u}_{0})$ from mode 0 and $\mathcal{D}(\mathcal{P}_{\ell}(\omega)\mathbf{u}_{\ell}+\mathcal{P}_{0}(\omega)\mathbf{u}_{0})-\mathcal{D}(\mathcal{P}_{0}(\omega)\mathbf{u}_{0})$ from modes $\ell=1,\dots,6$. The top row shows the input vorticity snapshots at (left panels) $Re=40$ with $D/D_{lam}=0.06,0.12,0.18,0.26$ from left to right, (right panels) $Re=100$ with $D/D_{lam}=0.016,0.032,0.047,0.064$ from left to right. Each column shows the contribution from mode $\ell$ to the reconstruction of the input vorticity snapshot. Each snapshot in a row shares the same colorbar to highlight the activations of different modes across dissipation rates.
@@
Figure 10: The contributions $\mathcal{D}(\mathcal{P}_{0}(\omega)\mathbf{u}_{0})$ from mode 0 and $\mathcal{D}(\mathcal{P}_{\ell}(\omega)\mathbf{u}_{\ell}+\mathcal{P}_{0}(\omega)\mathbf{u}_{0})-\mathcal{D}(\mathcal{P}_{0}(\omega)\mathbf{u}_{0})$ from modes $\ell=1,\dots,6$. The top row shows the input vorticity snapshots at (left panels) $Re=200$ with $D/D_{lam}=0.0064,0.0088,0.0102,0.0233$ from left to right, (right panels) $Re=400$ with $D/D_{lam}=0.0022,0.003,0.0035,0.0054$ from left to right. Each column shows the contribution from mode $\ell$ to the reconstruction of the input vorticity snapshot. Each snapshot in a row shares the same colorbar to highlight the activations of different modes across dissipation rates.
@@
To visualize the physical structures spanned by each mode, we decode the projections onto each mode back to physical space. We add the projection onto the $\ell=0$ mode when projecting onto $\ell>0$ modes as otherwise the output $\mathcal{D}(\mathcal{P}_{\ell}(\omega)\mathbf{u}_{\ell})$ is not physically realistic [for a similar discussion, see 34, 13]. Then, we subtract $\mathcal{D}(\mathcal{P}_{0}(\omega)\mathbf{u}_{0})$ from the decoded field, in an effort to isolate and interpret the contribution from the $\ell>0$ mode in physical space.
@@
The contributions $\mathcal{D}(\mathcal{P}_{0}(\omega)\mathbf{u}_{0})$ from mode 0 and $\mathcal{D}(\mathcal{P}_{\ell}(\omega)\mathbf{u}_{\ell}+\mathcal{P}_{0}(\omega)\mathbf{u}_{0})-\mathcal{D}(\mathcal{P}_{0}(\omega)\mathbf{u}_{0})$ from modes $\ell=1,\dots,6$ are shown in figure 9 for a selection of snapshots spanning the full range of dissipation rates at $Re=40$ and $Re=100$. The same modes for $Re=200$ and $Re=400$ are reported in figure 10. At all $Re$ considered, the $\ell=0$ mode bears some resemblance to the time-averaged vorticity state in the symmetry-reduced subspace across the full range of dissipation rates. This structure is also visually similar to many low-dissipation exact solutions in this flow (traveling waves and relative periodic orbits, see [37, 51, 34] – some examples at $Re=40$ are also included in an appendix). There is a weaker projection onto this state for the high dissipation snapshots at $Re=40$, whereas the projection at the higher $Re\in\{100,200,400\}$ is approximately equal across all snapshots presented. Similar to previous studies [34, 13], the $\ell=0$ mode is central to the latent representation – the singular value $\sigma_{0}$ is two orders of magnitude larger than $\sigma_{1}$ at $Re=40$. The $\ell=1$ mode also captures similar structures at both $Re=40$ and $Re=100$, and is dominated by paired diagonal elongated vortex cores of opposite circulation. This mode is visually similar to the most unstable eigenfunction of many low-dissipation exact solutions which resemble the mean flow state (see figure 21 in the appendix). Both $\ell=0$ and $\ell=1$ modes appear to be rotationally symmetric.
@@
One striking aspect of the decodes at all $Re$ is the sparsity in the representation – some modes are not active for the input snapshots. This is suggestive of a strong separation of modes for different dynamical processes, but is also likely due in part to the discrete symmetry reduction. The mapping to a fundamental chart is based on the first Fourier mode in the vertical. This can mean that snapshots on a trajectory can suddenly go through a ‘jump’ to another chart which suddenly shifts the local vortical features when symmetry reduction is performed (the first Fourier mode may not be apparent by eye, particularly in high dissipation snapshots), and the autoencoder therefore must learn to represent small scale vortices in different areas of the symmetry-reduced snapshots. This is perhaps apparent in the $\ell=3$ mode at $Re=400$ (figure 10) with a dipole in the upper corners of the decode, though in general the features across the modes reported in figures 9 and 10 are all physically distinct. In contrast to similar attempts to visualize physical representations of individual latent directions (e.g. in [13]) the decodes are all smooth and physically interpretable, even at high $Re$.
@@ V.2 Low-dimensional visualizations
Figure 11: Two-dimensional visualizations of the co-ordinate vector (12) at $Re=40$ using the t-SNE algorithm [52]. The $x-$ and $y-$axes represent the two degrees of freedom in the t-SNE visualization, which are determined such that the similarity between neighboring mapped data (measured by t-Student distributions) approximates the similarity in the original variables (measured by Gaussian distributions). Each snapshot is colored by a binned mean normalized dissipation rate $D/D_{lam}$ where $D_{lam}=Re/(2n^{2})$ is the laminar dissipation rate, using 25 bins in both $x-$ and $y-$directions. The dotted black lines indicate level curves of the Gaussian kernel density estimates of the representation, such that increasingly bold lines indicate increasing local density. The spanwise vorticity of representative states indicated by the corresponding marker visualize the distinct dynamical processes contained within the high density clusters of the visualization. The symbols and numbers in the high-dissipation regions are discussed later (see also figure 12).
@@
The sparseness of the representation in the latent SVD basis motivates a lower-dimensional visualization. We present two-dimensional representations of the modal projections of the embeddings (12) for each snapshot in the test dataset using the t-SNE algorithm [52] for $Re=40$, $Re=100$ and $Re=400$. When constructing these visualizations, we also include previously computed libraries of unstable periodic orbits [for full details see 14, 13, 38] and 10 short trajectories of 100 advective time units sampled at intervals of 0.1 time units, when generating the t-SNE visualization. These snapshots had to be included when running the t-SNE algorithm as it is a non-parametric self-supervised algorithm, but the UPOs and trajectory representations are only plotted in the low-dimensional visualizations when explicitly stated in the text. The two-dimensional visualization plane is divided into $25\times 25$ bins, and the mean dissipation rate of of the snapshots in each bin is computed. Each point in the representation represents a snapshot in the dataset, and is colored by this local mean dissipation rate. The contours in figures 11 and 14 indicate level curves of the local density of the visualization, computed via Gaussian kernels [53]. The qualitative features (the clusters) of the output are robust to different initializations of the algorithm.
@@
The two-dimensional visualizations in figures 11 and 14 should be compared to those computed previously in PHBK and [34] these were dominated by the learned shift-reflect symmetry, and the data appeared to form a large octagon with little other structure. The full symmetry reduction here, combined with the IRMAE approach, results in many more t-SNE clusters with distinct embedding features, each of which represents a distinctive ‘class’ of vorticity snapshots in physical space.
@@
At $Re=40$, representative flow states are sampled from a selection of low-dissipation t-SNE clusters and visualized in the right panels of figure 11, with snapshots identified on the t-SNE visualization with markers. The clusters all show the characteristic zig-zag “staircase” patterns in the vorticity, and are distinguished by the number and relative locations of dominant vortices. The detached clusters indicated by the light blue circle and salmon diamond are notably apart from the bulk of the visualization. They sit far from the higher dissipation events and contain the lowest dissipation fields with large patches of quiescent fluid. The clear clustering into many distinct classes of events – not visible in previous work [34, 13] – and apparent distinction of different types of high-dissipation dynamics motivates further investigation into the nature of bursting events and routes into/out of this region of state space.

### v2

@@ IV.2 Attractor dimension
Here we present the main result of this study – the scaling with $Re$ of the estimated dimension of the chaotic attractor. Implicit rank minimization drives the rank of the embedding space to an approximately minimal rank solution, which we use as (an upper bound for) our numerically estimated dimension. To estimate this minimal rank, we perform a singular value decomposition (SVD) on the latent data matrix $\mathbf{Z}\mathbf{Z}^{T}$ [27], where $\mathbf{Z}=[\dots z_{i}\dots]$ and $z_{i}=\mathscr{W}(\mathscr{E}(\omega_{i}))$ for all $\omega_{i}$ in the test dataset. The SVD of $\mathbf{Z}\mathbf{Z}^{T}=\mathbf{U}\bm{\Sigma}\mathbf{U}^{T}$ computes the singular vectors $\mathbf{u}_{j}\in\mathbb{R}^{d_{z}}$ where $\mathbf{U}=[\mathbf{u}_{0}\dots\mathbf{u}_{d_{z-1}}]$ (the left singular vectors of $\mathbf{Z}$) and singular values $\sigma_{j}\in\mathbb{R}$ where $\bm{\Sigma}=\text{diag}(\sigma_{0},\dots,\sigma_{d_{z-1}})$ (the square of the singular values of $\mathbf{Z}$). The number of non-negligible singular values in this spectrum is the estimate for $d_{\mathcal{A}}$ in the fully symmetry reduced subspace. Each of these singular vectors represents some important structure of the original data in the latent space, which will be investigated in more detail in §V.
@@
Figure 8: (Top) Singular values $\sigma_{i}$ in the singular value decompositions of the covariance of the latent data matrix at $Re=$ 25 (  ), 30 ( ), 35 ( ), 40 ( ), 60 ( ), 80 ( ), 100 ( ), 200 ( ) and 400 ( ). (Bottom) The estimated upper bound for the dimension of the turbulent attractor $\tilde{d}_{\mathcal{A}}$ against ReRe. The bold line shows the rigorous upper bound scaling $\sim Re^{4/3}(1+\log{(2^{5/2}\pi^{3}Re^{2})})^{1/3}$ and the dotted line shows the scaling $\sim Re^{1/3}$.
The singular values for each $Re$ considered are reported in the top panel of figure 8. In all cases there is a sharp drop before $\sigma_{j}$ becomes negligibly small at a critical index $j=c$ (defined as the first index where $\sigma_{c}<10^{-8}$), corresponding to the rank of the learned, fully symmetry-reduced, latent space. To compute the estimate for $d_{\mathcal{A}}$ in full state space, which we denote by $\tilde{d}_{\mathcal{A}}$, we add the continuous symmetry back into the estimate, such that $\tilde{d}_{\mathcal{A}}=c+1$ (reducing the discrete symmetries reduces the size of the state space, but not the dimension). We computed two estimates for $\tilde{d}_{\mathcal{A}}$ at each $Re$ by training two randomly initialized autoencoders from scratch. These estimates were found to be robust to the training – they differed only slightly between models at all $Re$. The smallest estimate at each $Re$ is reported in the bottom panel of figure 8.
The non-dimensional Grashof number $G$ in the rigorous upper bound [8, 9] of the global attractor dimension is related to the present definition of $Re$, with monochromatic $n=4$ Kolmogorov forcing, by $G=2^{5/2}\pi^{3}Re^{2}$. This upper bound then scales with $Re$ like $bRe^{4/3}(1+\log{(2^{5/2}\pi^{3}Re^{2})})^{1/3}$, and is indicated by the bold line in the bottom panel of figure 8 ($b$ is an unspecified constant). This bound and the lower bound in [10] for the global attractor both scale with the number of degrees of freedom in a two-dimensional domain ($Re^{4/3}$ is proportional to the square of the domain size $L$ over the Kraichnan scale) with a logarithmic correction.
In contrast, our data-driven estimate for the dimensionality of the chaotic attractor explored by long turbulent trajectories exhibits a much weaker scaling at high $Re$. While we observe some non-monotonic behavior in the bottom panel of figure 8 at lower $Re$, there is a sharp transition beyond $Re=40$ beyond which an approximate scaling $d_{\mathcal{A}}\sim Re^{1/3}$ is observed in the range $Re\in[60,400]$ – identified with the dotted line in the figure. Although this scaling was estimated over this limited range of $Re$ values, earlier work [38, 39] indicates that an ‘asymptotic’ regime (a clear power law scaling of dissipation with $Re$) is established for $Re\gtrsim 60$ and extends to at least $Re=1000$, which would suggest that our dimension scaling may well also continue to be robust beyond $Re=400$. The estimate $d_{\mathcal{A}}\sim Re^{1/3}$ is a weaker scaling than a previous result of $\sim Re^{0.78}$, which was computed via the Kaplan-Yorke conjecture [19]. However, this flow configuration was forced by a different profile and was damped at large scales, making a direct comparison challenging.
As a point of comparison for the non-linear dimension estimation by the IRMAE networks, we also performed a linear dimensionality reduction approach on the symmetry-reduced test datasets using principal component analysis (PCA) [54]. The singular values from a PCA performed at each $Re$ are reported in figure 9. There is no sharp cut-off in the singular values, especially at the higher $Re$, so it is not possible to meaningfully extract a dimension at each $Re$ using this approach.
Figure 9: Principal component analysis values $\sigma_{\ell}$ for the symmetry-reduced test datasets at $Re=$ 25 (  ), 30 ( ), 35 ( ), 40 ( ), 60 ( ), 80 ( ), 100 ( ), 200 ( ) and 400 ( ).
@@
One caveat with our estimates is that our models do not perfectly reconstruct their inputs, and this error does increase with $Re$. However, the markedly weaker scaling with $Re$ is consistent with the formation of coherent structures which are associated with a subset of wavenumber pathways in a Fourier representation. We also examined the relative error with respect to a reference trajectory as the model predictions are unrolled in time and found that the errors remain roughly constant for longer than a Lyapunov time (we compared at $Re=100$ where this time has been estimated at $T_{\lambda}=3.6$ by [55]). Another caveat is we also assume that we have sufficiently sampled the chaotic attractor with our test datasets of 50,000 snapshots at each $Re$. We now explore the representation of the turbulence learned in these models by decoding individual components from the SVD.
@@ V.1 Modal structures
We first examine the contribution to the vorticity field reconstruction from each of the singular vectors (modes) $\mathbf{u}_{\ell}$ in the SVD of $\mathbf{Z}\mathbf{Z}^{T}$. To do this, we project low and high dissipation snapshots onto the $\ell=0$ mode and combinations of this mode with higher modes. The projection operator onto mode $\ell$ in the embedding space for some vorticity field $\omega$ is defined by
@@
Figure 10: The contributions $\mathcal{D}(\mathcal{P}_{0}(\omega)\mathbf{u}_{0})$ from mode 0 and $\mathcal{D}(\mathcal{P}_{\ell}(\omega)\mathbf{u}_{\ell}+\mathcal{P}_{0}(\omega)\mathbf{u}_{0})-\mathcal{D}(\mathcal{P}_{0}(\omega)\mathbf{u}_{0})$ from modes $\ell=1,\dots,6$. The top row shows the input vorticity snapshots at (left panels) $Re=40$ with $D/D_{lam}=0.06,0.12,0.18,0.26$ from left to right, (right panels) $Re=100$ with $D/D_{lam}=0.016,0.032,0.047,0.064$ from left to right. Each column shows the contribution from mode $\ell$ to the reconstruction of the input vorticity snapshot. Each snapshot in a row shares the same colorbar to highlight the activations of different modes across dissipation rates. The bottom row shows the projections $\mathcal{P}_{\ell}$ (on a logarithmic scale) onto the first 25 latent modes for the snapshot in each column.
@@
Figure 11: The contributions $\mathcal{D}(\mathcal{P}_{0}(\omega)\mathbf{u}_{0})$ from mode 0 and $\mathcal{D}(\mathcal{P}_{\ell}(\omega)\mathbf{u}_{\ell}+\mathcal{P}_{0}(\omega)\mathbf{u}_{0})-\mathcal{D}(\mathcal{P}_{0}(\omega)\mathbf{u}_{0})$ from modes $\ell=1,\dots,6$. The top row shows the input vorticity snapshots at (left panels) $Re=200$ with $D/D_{lam}=0.0064,0.0088,0.0102,0.0233$ from left to right, (right panels) $Re=400$ with $D/D_{lam}=0.0022,0.003,0.0035,0.0054$ from left to right. Each column shows the contribution from mode $\ell$ to the reconstruction of the input vorticity snapshot. Each snapshot in a row shares the same colorbar to highlight the activations of different modes across dissipation rates. The bottom row shows the projections $\mathcal{P}_{\ell}$ (on a logarithmic scale) onto the first 25 latent modes for the snapshot in each column.
@@
To visualize the physical structures spanned by each mode, we decode the projections onto each mode back to physical space. We add the projection onto the $\ell=0$ mode when projecting onto $\ell>0$ modes i.e. $\mathcal{D}(\mathcal{P}_{\ell}(\omega)\mathbf{u}_{\ell}+\mathcal{P}_{0}(\omega)\mathbf{u}_{0})$ as otherwise the output $\mathcal{D}(\mathcal{P}_{\ell}(\omega)\mathbf{u}_{\ell})$ is not physically realistic [for a similar discussion, see 35, 14]. Then, we subtract $\mathcal{D}(\mathcal{P}_{0}(\omega)\mathbf{u}_{0})$ from the decoded field, in an effort to isolate and interpret the contribution from the $\ell>0$ mode in physical space.
@@
The contributions $\mathcal{D}(\mathcal{P}_{0}(\omega)\mathbf{u}_{0})$ from mode 0 and $\mathcal{D}(\mathcal{P}_{\ell}(\omega)\mathbf{u}_{\ell}+\mathcal{P}_{0}(\omega)\mathbf{u}_{0})-\mathcal{D}(\mathcal{P}_{0}(\omega)\mathbf{u}_{0})$ from modes $\ell=1,\dots,6$ are shown in figure 10 for a selection of snapshots spanning the full range of dissipation rates at $Re=40$ and $Re=100$. The same modes for $Re=200$ and $Re=400$ are reported in figure 11. At all $Re$ considered, the $\ell=0$ mode bears some resemblance to the time-averaged vorticity state in the symmetry-reduced subspace across the full range of dissipation rates. This structure is also visually similar to many low-dissipation exact solutions in this flow (traveling waves and relative periodic orbits, see [38, 56, 35] – some examples at $Re=40$ are also included in appendix B). There is a weaker projection onto this state for the high dissipation snapshots at $Re=40$, whereas the projection at the higher $Re\in\{100,200,400\}$ is approximately equal across all snapshots presented. Similar to previous studies [35, 14], the $\ell=0$ mode is central to the latent representation – the singular value $\sigma_{0}$ is two orders of magnitude larger than $\sigma_{1}$ at $Re=40$. The $\ell=1$ mode also captures similar structures at both $Re=40$ and $Re=100$, and is dominated by paired diagonal elongated vortex cores of opposite circulation. This mode is visually similar to the most unstable eigenfunction of many low-dissipation exact solutions which resemble the mean flow state (see figure 18 in appendix B).
@@
One striking aspect of the decodes at all $Re$ is the sparsity in the representation – some modes are not active for the input snapshots. For example, consider modes $\ell=1$ and $\ell=2$ in the left column of figure 10. The projections onto these modes $\mathcal{P}_{\ell}$ is plotted on a logarithmic scale in the bottom row of figure 10, demonstrating that $\mathcal{P}_{2}$ is two orders of magnitude smaller than $\mathcal{P}_{1}$. This is suggestive of a strong separation of modes for different dynamical processes, but is also likely due in part to the discrete symmetry reduction. The mapping to a fundamental chart is based on the first Fourier mode in the vertical. This can mean that snapshots on a trajectory can suddenly go through a ‘jump’ to another chart which suddenly shifts the local vortical features when symmetry reduction is performed (the first Fourier mode may not be apparent by eye, particularly in high dissipation snapshots), and the autoencoder therefore must learn to represent small scale vortices in different areas of the symmetry-reduced snapshots. This is perhaps apparent in the $\ell=3$ mode at $Re=400$ (figure 11) with a dipole in the upper corners of the decode, though in general the features across the modes reported in figures 10 and 11 are all physically distinct. In contrast to similar attempts to visualize physical representations of individual latent directions (e.g. in [14]) the decodes are all smooth and physically interpretable, even at high $Re$.
@@ V.2 Low-dimensional visualizations
Figure 12: Two-dimensional visualization of the co-ordinate vector (12) at $Re=40$ using the t-SNE algorithm [57]. (Left) The $x$- and $y$-axes represent the two degrees of freedom in the t-SNE visualization, which are determined such that the similarity between neighboring mapped data (measured by t-Student distributions) approximates the similarity in the original variables (measured by Gaussian distributions). Each snapshot is colored by a binned mean normalized dissipation rate $D/D_{lam}$ where $D_{lam}=Re/(2n^{2})$ is the laminar dissipation rate, using 25 bins in both $x$ and $y$-directions. The dotted black lines indicate level curves of the Gaussian kernel density estimates of the representation, such that increasingly bold lines indicate increasing local density. The spanwise vorticity of representative states indicated by the arrows visualize the distinct dynamical processes contained within the high density clusters of the visualization. Snapshots are extracted belonging to the bursting event class $i=1,2$, and the average projection onto each mode $\langle\mathcal{P}_{\ell}\rangle_{i}$ of class $i$ is computed. (Right top) The difference $\Delta\langle\mathcal{P}_{\ell}\rangle\coloneq\langle\mathcal{P}_{\ell}\rangle_{1}-\langle\mathcal{P}_{\ell}\rangle_{2}$ is shown for each mode $\ell$. A subset of modes $\ell\in L_{d}$ (red bars) appears to be mostly responsible for delineating the two bursting event classes. (Right bottom) Two representative snapshots $\omega$ indicated by fuchsia (top panels) and cyan (bottom panels) markers from the two classes of bursting events are shown. The decodes $\mathcal{D}(\sum_{\ell\in L_{d}}\mathcal{P}_{\ell}(\omega)u_{\ell})$ are shown directly to the right of each representative snapshot $\omega$.
@@
The sparseness of the representation in the latent SVD basis motivates a lower-dimensional visualization. We present two-dimensional representations of the modal projections of the embeddings (12) for each snapshot in the test dataset using the t-SNE algorithm [57] for $Re=40$, $Re=100$ and $Re=400$. As local pairwise distances between snapshots in the (truncated) latent space are ‘preserved’ in the t-SNE visualizations [57], nearby snapshots in these visualizations reflect nearby snapshots in the truncated latent space. When constructing these visualizations, we also include previously computed libraries of unstable periodic orbits [for full details see 15, 14, 39] and 10 short trajectories of 100 advective time units sampled at intervals of 0.1 time units, when generating the t-SNE visualization. These snapshots had to be included when running the t-SNE algorithm as it is a non-parametric self-supervised algorithm, but the UPOs and trajectory representations are only plotted in the low-dimensional visualizations when explicitly stated in the text. The two-dimensional visualization plane is divided into $25\times 25$ bins, and the mean dissipation rate of of the snapshots in each bin is computed. Each point in the representation represents a snapshot in the dataset, and is colored by this local mean dissipation rate. The contours in figures 12 and 14 indicate level curves of the local density of the visualization, computed via Gaussian kernels [58]. The qualitative features (the clusters) of the output are robust to different initializations of the algorithm.
@@
The two-dimensional visualizations in figures 12 and 14 should be compared to those computed previously in PHBK and [35] – these were dominated by the learned shift-reflect symmetry, and the data appeared to form a large octagon with little other structure. The full symmetry reduction here, combined with the IRMAE approach, results in many more t-SNE clusters with distinct embedding features, each of which represents a distinctive ‘class’ of vorticity snapshots in physical space.
@@
At $Re=40$, representative flow states are visualized in figure 12 from a selection of low-dissipation t-SNE clusters, with their respective positions in the t-SNE visualization indicated by arrows. The clusters all show the characteristic zig-zag “staircase” patterns in the vorticity, and are distinguished by the number and relative locations of dominant vortices. Two clusters on the left of the t-SNE visualization are notably detached from the bulk of the visualization. They sit far from the higher dissipation events and contain the lowest dissipation fields with large patches of quiescent fluid (see the representative snapshots with a white circle in the top left corner). The clear clustering into many distinct classes of events – not visible in previous work [35, 14] – and apparent distinction of different types of high-dissipation dynamics motivates further investigation into the nature of bursting events and routes into/out of this region of state space.
V.2.1 Bursting classes